Calculate the mole fraction of 25% aqueous hydrochloric acid solution


1
Expert's answer
2022-04-17T11:04:29-0400

25% solution mean that "\frac{m(HCl)}{m(HCl)+m(H_2O)}=0.25"

Let m(HCl)+m(H2O)=100 g, then

m(HCl)=25 g

m(H2O)=100-25=75 g

M(HCl)=1+35.5=36.5 g/mol

M(H2O)=1x2+16=18 g/mol

n(HCl)=25/36.5=0.68 mol

n(H2O)=75/18=4.2 mol

"w(HCl)=0.68\/4.2=0.162"